### v1

# Cohomology of Tanabe algebras

Abstract.
In this paper we study the (co)homology of Tanabe algebras, which are a family of subalgebras of the partition algebras exhibiting a Schur-Weyl duality with certain complex reflection groups. The homology of the partition algebras has been shown to be related to the homology of the symmetric groups by Boyd-Hepworth-Patzt and the results they obtain depend on a parameter. In all known results, the homology of a diagram algebra is dependent on one of two things: the invertibility of a parameter in the ground ring or the parity of the positive integer indexing the number of pairs of vertices. We show that the (co)homology of Tanabe algebras is isomorphic to the (co)homology of the symmetric groups and that this is independent of both the parameter and the parity of the index. To the best of our knowledge, this is the first example of a result of this sort. Along the way we will also study the (co)homology of the uniform block permutation algebra and the totally propagating partition algebra as well collecting cohomological analogues of known results for the homology of partition algebras and Jones annular algebras.
Key words and phrases:
diagram algebras, cohomology of algebras, partition algebras, Tanabe algebras
1991 Mathematics Subject Classification:
16E40, 20J06, 16E30
1. Introduction
The homology of diagram algebras is an emerging area of study. Diagram algebras are examples of augmented algebras and their homology and cohomology can be defined in terms of certain $\operatorname{Tor}$ and $\operatorname{Ext}$ groups following [Ben91, Definition 2.4.4]. Examples of algebras that have been studied include the Temperley-Lieb algebras ([BH24]), the Brauer algebras ([BHP21]), the partition algebras ([BHP23, Boy24b]), the Jones annular algebras ([Boy24b]), the rook algebras ([Boy24a]), and the rook-Brauer algebras ([Boy24a]).
The partition algebras, $P_{n}(\delta)$ (where $n$ is a positive integer and $\delta$ is a parameter in a unital, commutative ground ring), were introduced independently by Martin ([Mar94]) and Jones ([Jon94a]) to study the Potts model. Jones showed that the partition algebras exhibit a Schur-Weyl duality with the symmetric groups. Tanabe ([Tan97]) introduced a family of subalgebras, $\mathcal{T}_{n}(\delta,r)$, of the partition algebra, one for each positive integer $r$, and demonstrated that these subalgebras exhibit a Schur-Weyl duality with certain complex reflection groups. Our treatment of the Tanabe algebras will follow that of Orellana ([Ore07]).
In this paper we will study the (co)homology of the Tanabe algebras. The homology of partition algebras has been studied by Boyd, Hepworth and Patzt ([BHP23]). Their results split into two cases, which depend on the parameter $\delta$. They show that if the parameter $\delta$ is invertible in the ground ring, then the homology of the partition algebras is globally isomorphic to the homology of the symmetric groups. However, if $\delta$ is not invertible then the homology of the partition algebras is only isomorphic to the homology of the symmetric groups in a range. This is a recurring phenomenon in this very young field. The homology of the Temperley-Lieb algebras ([BH24]) and the homology of the Brauer algebras ([BHP21]) both exhibit different behaviour depending on the parameter. The homology of some diagram algebras also exhibit different behaviours depending on the parity of the index $n$, which counts the number of pairs of vertices in a basis element. Such results exist for the homology of Temperley-Lieb algebras ([Sro22, Theorem A] and [Boy24a, Theorem 1.2]) and the Brauer algebras ([Boy24a, Theorem 1.3]).
In this paper we will show that for $r\geqslant 2$ the (co)homology of the Tanabe algebras, $\mathcal{T}_{n}(\delta,r)$, is isomorphic to the (co)homology of the symmetric group $\Sigma_{n}$ and that this is independent of both the parameter $\delta$ and the index $n$. To the best of our knowledge this is the first example of a result of this type to appear in the literature. It appears in the text as Part (1) of Theorem 6.0.3.
Theorem.
Let $\delta$ be any element of a unital, commutative ground ring $k$. Let $r\geqslant 2$. There exist isomorphisms of graded $k$-modules
$\operatorname{Tor}_{\star}^{\mathcal{T}_{n}(\delta,r)}\left(\mathbbm{1},\mathbbm{1}\right)\cong H_{\star}(\Sigma_{n},\mathbbm{1})\quad\text{and}\quad\operatorname{Ext}_{\mathcal{T}_{n}(\delta,r)}^{\star}\left(\mathbbm{1},\mathbbm{1}\right)\cong H^{\star}(\Sigma_{n},\mathbbm{1}).$
Along the way, we will show that the same methods can be used to show that the (co)homology of two other families of subalgebras of the partition algebra, called the totally propagating partition algebras and the uniform block permutation algebras, can also be identified with the (co)homology of the symmetric groups.
We will use techniques which were developed in [Boy24b] to study the homology of partition algebras and Jones annular algebras. These techniques, themselves inspired by work of Sroka ([Sro22]), involve a chain complex called the Mayer-Vietoris complex. In the case of the partition algebras, Boyde shows that this is a partial resolution of the trivial module to obtain isomorphisms between the homology of the partition algebras and the homology of the symmetric groups. However, in the case of the Tanabe algebras (and also the totally-propagating partition algebras and uniform block permutation algebras), the Mayer-Vietoris complex turns out to be a genuine resolution. The key reason for this is that the basis diagrams of these algebras are not allowed to have isolated vertices.
These techniques will also allow us to deduce cohomological versions of Boyde’s results for the Jones annular algebras and the partition algebras. As a corollary, we prove a cohomological stability result for the partition algebras after the fashion of [BHP23, Theorem B].
The paper is structured as follows. In Section 2, we recall the definitions of partition algebras, Tanabe algebras, totally propagating partition algebras and uniform block permutation algebras. In Section 3 we recall the definition of the (co)homology of an augmented algebra. In Section 4 we will recall the necessary technical results from [Boy24b] as well as proving the cohomological analogue of [Boy24b, Theorem 1.7]. We will use this to deduce cohomological versions of Boyde’s results for the Jones annular algebras. In Section 5 we briefly spell out the cohomological analogue of [Boy24b, Theorem 1.1], demonstrating that the cohomology of the partition algebras is stably isomorphic to the cohomology of the symmetric groups. We deduce that the partition algebras exhibit cohomological stability as a corollary. We also adapt a method of Boyd, Hepworth and Patzt to show that the cohomology of partition algebras with invertible parameter is globally isomorphic to the cohomology of the symmetric groups. In Section 6 we prove our main result, showing that the (co)homology of the Tanabe algebras is globally isomorphic to the (co)homology of the symmetric groups, without any conditions on the parameter or the parity of the index. We also identify the (co)homology of the totally propagating partition algebras and the uniform block permutation algebras.
Acknowledgements
We would like to thank James Brotherston and Natasha Cowley for helpful and interesting conversations whilst writing this paper. We would like to thank James Cranch and Sarah Whitehouse for the their feedback and support in this project and related work. We’d like to thank Rachael Boyd and Richard Hepworth for interesting conversations at the 2024 British Topology Meeting in Aberdeen. We would like to thank Guy Boyde for his comments and feedback on a previous draft of this paper.
1.1. Conventions
Throughout, unless otherwise stated, $k$ will be a unital, commutative ring. The notation $\underline{n}$ will be used for the set $\left\{1,\dotsc,n\right\}$.
2. An aggregation of algebras
In this section we collect the definitions of the algebras that we will consider in this paper, namely the partition algebras, the Tanabe algebras, the totally propagating partition algebras and the uniform block permutation algebras.
2.1. Partition algebras
Definition 2.1.1.
A partition $n$-diagram is a graph on two columns of $n$ vertices where each edge is incident to two distinct vertices and there is at most one edge between any two vertices. By convention, the vertices down the left-hand column will be labelled by $1,\dots,n$ in ascending order from top to bottom and the vertices down the right-hand column will be labelled by $\overline{1},\dotsc,\overline{n}$ in ascending order from top to bottom.
These diagrams are called partition $n$-diagrams because the connected components of the graph determine a partition of the set $\left\{1,\overline{1},\dotsc,n,\overline{n}\right\}$. We say two partition $n$-diagrams are equivalent if they determine the same partition of the set $\left\{1,\overline{1},\dotsc,n,\overline{n}\right\}$. Henceforth, when referring to a diagram, we will mean its equivalence class.
Definition 2.1.2.
We collect some important terminology for graphs that will be used throughout the rest of the paper.
  1. (1)
    An edge that connects the left-hand column of vertices to the right-hand column of vertices will be called a propagating edge.
  2. (2)
    A connected component which contains vertices in both columns (that is, a connected component containing a propagating edge) will be called a propagating component.
  3. (3)
    An edge that connects two vertices in the same column will be called a non-propagating edge.
  4. (4)
    A vertex not connected to any other by an edge will be called an isolated vertex.
  5. (5)
    We will refer to the number of vertices in a connected component of a partition $n$-diagram as the cardinality of the connected component.
  6. (6)
    Any diagram having precisely $n$ propagating components will be called an permutation diagram. All other diagrams will be referred to as non-permutation diagrams.
Definition 2.1.3.
Let $\delta\in k$. The partition algebra, $P_{n}(\delta)$, is the $k$-algebra with basis consisting of all partition $n$-diagrams with the multiplication defined by the $k$-linear extension of the following product of diagrams. Let $d_{1}$ and $d_{2}$ be partition $n$-diagrams. The product $d_{1}d_{2}$ is obtained by the following procedure:
  • Place the diagram $d_{2}$ to the right of the diagram $d_{1}$ and identify the vertices $\overline{1},\dotsc,\overline{n}$ in $d_{1}$ with the vertices $1,\dotsc,n$ in $d_{2}$. Call this diagram with three columns of vertices $d_{1}\ast d_{2}$. We drop the labels of the vertices in the middle column and we preserve the labels of the left-hand column and right-hand column.
  • Count the number of connected components that lie entirely within the middle column. Call this number $\alpha$.
  • Make a new partition $n$-diagram, $d_{3}$, as follows. Given distinct vertices $x$ and $y$ in the set $\left\{1,\overline{1},\dotsc,n,\overline{n}\right\}$, $d_{3}$ has an edge between $x$ and $y$ if there is a path from $x$ to $y$ in $d_{1}\ast d_{2}$.
  • We define the composite $d_{1}d_{2}=\delta^{\alpha}d_{3}$.
We note that this product is associative and well-defined up to equivalence of partition $n$-diagrams. The identity element consists of the diagram with $n$ horizontal edges.
Example 2.1.4.
Here is an example of the composition defined above. Suppose we have diagrams
$d_{1}$$=$$\overline{4}$$\overline{3}$$\overline{2}$$\overline{1}$$d_{2}$$=$$\overline{4}$$\overline{3}$$\overline{2}$$\overline{1}$
in $P_{n}(\delta)$. In this case we have
[…]
We refer the reader to [Jon94a], [Blo03] and [BHP23] for some more examples of composing partition diagrams. (We note that some authors work with two rows of vertices rather than two columns and compose diagrams vertically rather than horizontally.)
2.2. Tanabe algebras
In this subsection we define the Tanabe algebras. If we fix positive integers $r$ and $n$, the Tanabe algebra is defined as the subalgebra of the partition algebra spanned $k$-linearly by those partition $n$-diagrams such that, for each connected component, the difference between the number of vertices in the left and right columns is congruent to zero modulo $r$.
Definition 2.2.1.
Let $d$ be a partition $n$-diagram. Let $C$ be a connected component of $d$, let $L(C)$ denote the number of vertices in $C$ in the left-hand column of $d$ and let $R(C)$ denote the number of vertices in $C$ in the right-hand column of $d$. We define
$\kappa(C)=\left\lvert L(C)-R(C)\right\rvert.$
Definition 2.2.2.
Let $\delta\in k$. Let $r\geqslant 1$. The Tanabe algebra $\mathcal{T}_{n}(\delta,r)$ is the subalgebra of $P_{n}(\delta)$ spanned $k$-linearly by the partition $n$-diagrams such that $\kappa(C)\equiv 0\,(\text{mod}\,r)$ for each connected component $C$.
Remark 2.2.3.
If we take $r=1$, we recover the partition algebra, $P_{n}(\delta)$. When $r=2$ the family of algebras are sometimes called the even partition algebras or the parity matching algebras (see [Scr24] for instance).
2.3. Totally propagating partition algebras
The totally propagating partition algebras have been studied in [KM08], where they are shown to be the Schur-Weyl duals of the rook algebras (see [HR01, Sol02, Gro02, Hal04, Xia16]). The underlying monoid had appeared in earlier papers: [FL98, Mal07, EEF08] (under the name dual symmetric inverse monoid). The totally propagating partition algebras have been further studied in [MS21].
Definition 2.3.1.
Recall that a connected component of a partition $n$-diagram is called propagating if it contains vertices in both the left-hand column and the right-hand column. A partition $n$-diagram is called totally propagating if all of its connected components are propagating.
Definition 2.3.2.
The totally propagating partition algebra, $TPP_{n}$, is the subalgebra of $P_{n}(\delta)$ spanned $k$-linearly by the totally propagating partition $n$-diagrams. We note that in the procedure for composing two totally propagating partition $n$-diagrams $d_{1}$ and $d_{2}$, the diagram $d_{1}\ast d_{2}$ can have no connected components that lie entirely within the middle column so we drop the parameter $\delta$ from the notation.
2.4. Uniform block permutation algebras
The uniform block permutation algebras were first introduced by Kosuda, under the name party algebras [Kos00] (see also [Kos06]). The name uniform block permutation algebra was coined by FitzGerald [Fit03]. See [AO08, OSSZ22, OSSZ24] for more recent results about the uniform block permutation algebras.
Definition 2.4.1.
Recall the notation of Definition 2.2.1. Let $d$ be a partition $n$-diagram. A connected component $C$ of $d$ is called uniform if $L(C)=R(C)$, that is, if $C$ has an equal number of vertices in both the left-hand and right-hand column of $d$. The diagram $d$ is called uniform if every connected component of $d$ is uniform. Some authors refer to such a partition as being balanced (see [Har18] for instance).
Definition 2.4.2.
The uniform block permutation algebra, $U_{n}$, is the subalgebra of $P_{n}(\delta)$ spanned $k$-linearly by the uniform partition $n$-diagrams. We note that in the procedure for composing two uniform partition $n$-diagrams $d_{1}$ and $d_{2}$, the diagram $d_{1}\ast d_{2}$ can have no connected components that lie entirely within the middle column so we drop the parameter $\delta$ from the notation.
3. Cohomology of augmented algebras
Recall that a $k$-algebra is said to be augmented if it comes equipped with a $k$-algebra map $\varepsilon\colon A\rightarrow k$, which is called the augmentation.
Each of the algebras in Section 2 can be equipped with an augmentation. We define these augmentations as follows. For any algebra $A$ defined in 2, let $I_{n-1}$ denote the two-sided ideal in $A$ spanned $k$-linearly by all diagrams having fewer than $n$ propagating components (one can see that this is a two-sided ideal by checking that the composite of two diagrams with $i$ propagating components and $j$ propagating components respectively is a scalar multiple of a diagram with at most $\min(i,j)$ propagating components). Since $I_{n-1}$ is a two-sided ideal, we have a well-defined augmentation, which we give in the following definition.
Definition 3.0.1.
Let $A$ be any algebra defined in Section 2. We equip $A$ with the augmentation $\varepsilon\colon A\rightarrow k$ that sends the permutation diagrams to $1\in k$ and all non-permutation diagrams to $0\in k$.
We note that $I_{n-1}$ is not the augmentation ideal, since the augmentation ideal also contains certain formal $k$-linear combinations of permutation diagrams.
Definition 3.0.2.
For any algebra, $A$, defined in Section 2, the trivial module, $\mathbbm{1}$, consists of a single copy of the ground ring $k$, where $A$ acts on $k$ via the augmentation.
We now recall the definition of the homology and cohomology of an augmented algebra, from [Ben91, Definition 2.4.4].
Definition 3.0.3.
Let $A$ be an augmented $k$-algebra and let $\mathbbm{1}$ denote a copy of the ground ring $k$, where $A$ acts on $\mathbbm{1}$ via the augmentation map. We define the homology and cohomology of $A$ to be
$\operatorname{Tor}_{\star}^{A}(\mathbbm{1},\mathbbm{1})\quad\text{and}\quad\operatorname{Ext}_{A}^{\star}(\mathbbm{1},\mathbbm{1})$
respectively.
4. Results on the cohomology of algebras
We recall the definitions of $k$-free idempotent left cover and the Mayer-Vietoris complex from [Boy24b] and prove the cohomological analogue of [Boy24b, Theorem 1.7].
4.1. Idempotent left covers and the Mayer-Vietoris complex
The material in this subsection comes from [Boy24b, Sections 1 and 2].
Definition 4.1.1.
Let $A$ be a $k$-algebra. Let $I$ be a two-sided ideal of $A$. Let $w\geqslant h\geqslant 1$. An idempotent left cover of $I$ of height $h$ and width $w$ is a collection of left ideals $J_{1},\dotsc,J_{w}$ in $A$ such that
  • $J_{1}+\cdots+J_{w}=I$;
  • for $S\subset\underline{w}$ with $\left\lvert S\right\rvert\leqslant h$, the intersection
    $\bigcap_{i\in S}J_{i}$
    is either zero or is a principal left ideal generated by an idempotent.
If $I$ is free as a $k$-module, then an idempotent left cover is said to be $k$-free if there is a choice of $k$-basis for $I$ such that each $J_{i}$ is free on a subset of this basis.
Definition 4.1.2.
Let $A$ be a $k$-algebra. Let $I\subset A$ be a two-sided ideal. Let $J_{1},\dotsc,J_{w}$ be an idempotent left cover of $I$. The Mayer-Vietoris complex associated to the idempotent left cover, $C_{\star}$, is the chain complex of left $A$-modules defined as follows. We set
$C_{p}=\underset{\left\lvert S\right\rvert=p}{\bigoplus_{S\subset\underline{w}}}\bigcap_{i\in S}J_{i}$
for $1\leqslant p\leqslant w$. We set $C_{0}=A$, $C_{-1}=A/I$ and $C_{n}=0$ for $n>w$ and $n<-1$.
The differential $C_{0}\rightarrow C_{-1}$ is the projection map $A\rightarrow A/I$. The differential $C_{1}\rightarrow C_{0}$ is the direct sum of the inclusion of the left ideals $J_{i}\rightarrow A$. For $p\geqslant 2$, the differential $C_{p}\rightarrow C_{p-1}$ is defined on the summand $\cap_{i\in S}J_{i}$ by
$x\mapsto\sum_{j\in S}(-1)^{\#(S,j)}i_{(S,j)}(x)$
where #$\#(S,j)$ is the number of elements of $S$ that are less than $j$ and $i_{(S,j)}$ is the inclusion
$\bigcap_{i\in S}J_{i}\rightarrow\bigcap_{i\in S\setminus\left\{j\right\}}J_{i}.$
Recall that for a left $A$-module $M$, a partial projective resolution of length $h$ of $M$ by left $A$-modules is an exact sequence
$P_{h}\rightarrow P_{h-1}\rightarrow\cdots\rightarrow P_{0}\rightarrow M\rightarrow 0$
where each $P_{i}$ is a projective left $A$-module.
The Mayer-Vietoris complex satisfies the following important property ([Boy24b, Proposition 2.5]).
Proposition 4.1.3.
Let $A$ be a $k$-algebra. Let $I$ be a two-sided ideal of $A$. Let $J_{1},\dotsc,J_{w}$ be a $k$-free idempotent left cover of $I$ of height $h$.
The truncation, $C_{\star}^{\leqslant h}$ of the Mayer-Vietoris complex associated to the idempotent left cover is a length $h$ partial projective resolution of $A/I$ by left $A$-modules with $C_{0}=A$. The partial projective resolution has the additional property that $X\otimes_{A}C_{p}^{\leqslant h}=0$ for $p\geqslant 1$ for any right $A$-module $X$ on which $I$ acts trivially. If $h=w$ then $C_{\star}^{\leqslant h}=C_{\star}$ is a projective resolution of $A/I$ by left $A$-modules.
4.2. A stable isomorphism on cohomology
The following is the cohomological version of [Boy24b, Theorem 1.7].
Proposition 4.2.1.
Let $M$ and $N$ be right $A$-modules. Let $I$ be a two-sided ideal that acts trivially on $M$ and $N$.
Suppose that there exists a partial projective resolution of length $h$, $C_{\star}^{\leqslant h}$, of $A/I$ by left $A$-modules such that $C_{0}^{\leqslant h}=A$ and such that $X\otimes_{A}C_{p}^{\leqslant h}=0$ for $p\geqslant 1$ for any right $A$-module $X$ on which $I$ acts trivially.
There is a natural isomorphism of $k$-modules
$\operatorname{Ext}_{A}^{q}(M,N)\cong\operatorname{Ext}_{A/I}^{q}(M,N)$
for $q\leqslant h$.
Proof.
When $q\leqslant h$, we will show that $\operatorname{Ext}_{A}^{q}(M,N)$ and $\operatorname{Ext}_{A/I}^{q}(M,N)$ are the cohomology of the same cochain complex.
Let $F_{\star}$ be a free resolution of $M$ by right $A$-modules. We know that $\operatorname{Ext}_{A}^{\star}(M,N)$ is the cohomology of the cochain complex $\operatorname{Hom}_{A}(F_{\star},N)$.
Since $I$ acts trivially on $N$, we have an isomorphism of cochain complexes
$\operatorname{Hom}_{A}(F_{\star},N)\cong\operatorname{Hom}_{A/I}\left(F_{\star}\otimes_{A}(A/I),N\right)$
by extension and restriction of scalars.
The same argument applied in [Boy24b, Theorem 2.7] (which uses the complex $C_{\star}^{\leqslant h}$ and the fact that $I$ acts trivially on $M$) shows that $F_{\star}^{\leqslant h+1}\otimes_{A}(A/I)$ is a length $h+1$ partial projective resolution of $M$ by right $A/I$-modules. Therefore $\operatorname{Hom}_{A/I}\left(F_{\star}\otimes_{A}(A/I),N\right)$ computes $\operatorname{Ext}_{A/I}(M,N)$ in degrees $q\leqslant h$ as required. ∎
Proposition 4.2.2.
Let $A$ be an augmented $k$-algebra with trivial module $\mathbbm{1}$. Let $I$ be a two-sided ideal of $A$ which is free as a $k$-module and which acts as multiplication by $0\in k$ on $\mathbbm{1}$. Suppose that there exists a $k$-free idempotent left cover of $I$ of height $h$ and width $w$. There is a natural isomorphism of $k$-modules
$\operatorname{Ext}_{A}^{q}(\mathbbm{1},\mathbbm{1})\cong\operatorname{Ext}_{A/I}^{q}(\mathbbm{1},\mathbbm{1})$
for $q\leqslant h$. Furthermore, if $h=w$, then the isomorphism holds for all $q$.
Proof.
This follows from Proposition 4.2.1 and Proposition 4.1.3 ([Boy24b, Proposition 2.4]), taking $M=N=\mathbbm{1}$. ∎
Boyde uses his Theorem 1.7 to prove results about the homology of Jones annular algebras, $J_{n}(\delta)$. Loosely speaking, the Jones annular algebras can be defined as the $k$-linear span of partition $n$-diagrams such that each connected component has size two and the diagram can be represented as a planar graph on an annulus (see [Jon94b, GL96, Boy24b] for precise definitions). The Jones annular algebras are augmented by sending the diagrams that represent cyclic permutations to $1\in k$ and all other diagrams to $0\in k$. We note here that Proposition 4.2.2 applies with the $k$-free idempotent left cover constructed in [Boy24b, Section 4]. We deduce the following result.
Theorem 4.2.3.
Let $\delta\in k$. Let $C_{n}$ denote the cyclic group of order $n$. There is a natural isomorphism
$\operatorname{Ext}_{J_{n}(\delta)}^{q}\left(\mathbbm{1},\mathbbm{1}\right)\cong H^{q}(C_{n},\mathbbm{1})$
for $q\leqslant\frac{n}{2}-1$.
Furthermore, if $n$ is odd or if $\delta$ is invertible, the isomorphism holds for all $q$.
5. Cohomology of partition algebras
In this section we show that for any $\delta\in k$, the cohomology of the partition algebras is stably isomorphic to the group cohomology of the symmetric groups using Proposition 4.2.2. As a corollary, we deduce that the partition algebras satisfy cohomological stability. Furthermore we show that if $\delta$ is invertible, then the cohomology of the partition algebras is globally isomorphic to the group cohomology of the symmetric groups, using methods introduced in [BHP23].
5.1. Stable isomorphism
We begin by showing that the cohomology of the partition algebras is stably isomorphic to the cohomology of the symmetric groups.
Theorem 5.1.1.
There is a natural isomorphism of $k$-modules
$\operatorname{Ext}_{P_{n}(\delta)}^{q}(\mathbbm{1},\mathbbm{1})\cong H^{q}(\Sigma_{n},\mathbbm{1})$
for $q\leqslant n-1$.
Proof.
Let $A=P_{n}(\delta)$. Recall that $I_{n-1}$ is the two-sided ideal of $A$ spanned $k$-linearly by the non-permutation diagrams. In particular, there is an isomorphism of $k$-algebras $P_{n}(\delta)/I_{n-1}\cong k[\Sigma_{n}]$. In order to prove the theorem we make use of Boyde’s $k$-free idempotent left cover of $I_{n-1}$.
For $i\in\underline{n}$, let $K_{i}$ denote the left ideal in $P_{n}(\delta)$ spanned $k$-linearly by the diagrams where the vertex $\overline{i}$ is an isolated vertex. For distinct $i$ and $j$ in $\underline{n}$ with $i<j$, we let $L_{i,j}$ denote the left ideal in $P_{n}(\delta)$ spanned $k$-linearly by the diagrams where $\overline{i}$ and $\overline{j}$ are in the same connected component.
Boyde ([Boy24b, Section 3]) shows that the left ideals $K_{i}$ and $L_{i,j}$ form a $k$-free idempotent left cover of $I_{n-1}$ of height $n-1$. The theorem now follows from Proposition 4.2.2. ∎
As a corollary, we can deduce the cohomological analogue of [BHP23, Theorem B].
Corollary 5.1.2.
The inclusion map $P_{n-1}(\delta)\rightarrow P_{n}(\delta)$ induces a map on cohomology
$\operatorname{Ext}_{P_{n}(\delta)}^{q}(\mathbbm{1},\mathbbm{1})\rightarrow\operatorname{Ext}_{P_{n-1}(\delta)}^{q}(\mathbbm{1},\mathbbm{1})$
which is an isomorphism in degrees $n\geqslant 2q+1$. Furthermore,
$\underset{n\rightarrow\infty}{\lim}\,\operatorname{Ext}_{P_{n}(\delta)}^{\star}(\mathbbm{1},\mathbbm{1})\cong\underset{n\rightarrow\infty}{\lim}\,H^{\star}\left(\Sigma_{n},\mathbbm{1}\right).$
Proof.
This follows from Theorem 5.1.1 and Nakaoka’s result on the cohomology stability of the symmetric groups [Nak60]. ∎
5.2. Invertible parameter
In this subsection we will prove that if the parameter $\delta\in k$ is invertible then the cohomology of $P_{n}(\delta)$ is globally isomorphic to the cohomology of $\Sigma_{n}$. We do this by finding a replacement of the cohomological version of Shapiro’s lemma for the partition algebras. This should be seen as the cohomological companion to [BHP23, Theorem 5.1].
Definition 5.2.1.
Let $J_{m}$ denote the left ideal of $P_{n}(\delta)$ consisting of all diagrams in which, amongst the vertices labelled $\left\{\overline{n-m+1},\dotsc,\overline{n}\right\}$, there is at least one singleton, or one pair of vertices in the same connected component.
Remark 5.2.2.
As noted in [BHP23, Lemma 5.2], for $m\leqslant n$, $P_{n}(\delta)$ is a right $k[\Sigma_{m}]$-module and $P_{n}(\delta)/J_{m}$ is a free right $k[\Sigma_{m}]$-module.
In the following two results, we will drop the $\delta$ and denote the partition algebra simply by $P_{n}$ in order to ease notation.
Lemma 5.2.3.
For $m\leqslant n$, there is an isomorphism of right $P_{n}$-modules
$\operatorname{Hom}_{k[\Sigma_{m}]}\left(P_{n}/J_{m},\mathbbm{1}\right)\cong\operatorname{Hom}_{P_{m}}\left(P_{n},\mathbbm{1}\right)$
given by the maps
$\varphi\colon\operatorname{Hom}_{k[\Sigma_{m}]}\left(P_{n}/J_{m},\mathbbm{1}\right)\rightarrow\operatorname{Hom}_{P_{m}}\left(P_{n},\mathbbm{1}\right)$
determined by $\varphi(f)(x)=f([x])$ and
$\psi\colon\operatorname{Hom}_{P_{m}}\left(P_{n},\mathbbm{1}\right)\rightarrow\operatorname{Hom}_{k[\Sigma_{m}]}\left(P_{n}/J_{m},\mathbbm{1}\right)$
determined by $\psi(g)([x])=g(x)$.
Proof.
It suffices to show that the maps are well-defined as it is then clear that they are mutually inverse.
We follow the conventions of [BHP23, Lemma 5.3]. We regard $J_{m}$ as an ideal of $P_{n}(\delta)$ and write $J_{m}\cap P_{m}$ for the corresponding ideal in $P_{m}(\delta)$. As noted in [BHP23, Lemma 5.3], $P_{m}\cong k[\Sigma_{m}]\oplus\left(J_{m}\cap P_{m}\right)$. In particular, for $\sigma\in J_{m}$, we can write $\sigma$ as a sum of terms of the form $x\sigma^{\prime}$ with $x\in P_{n}$ and $\sigma^{\prime}\in J_{m}$.
To show that the map $\varphi$ is well-defined, we must show that $\left(\varphi(f)(x)\right)\sigma=\varphi(f)(x\sigma)$ for $\sigma\in P_{m}$. By the above, it suffices to show this for $\sigma\in\Sigma_{m}$ and for $\sigma\in J_{m}\cap P_{m}$.
For $\sigma\in\Sigma_{m}$, we have $\left(\varphi(f)(x)\right)\sigma=\varphi(f)(x)=f([x])$ since $\sigma$ acts as multiplication by $1\in k$ on $\mathbbm{1}$. On the other hand, $\varphi(f)(x\sigma)=f([x\sigma])=f([x])\sigma=f([x])$, where the penultimate equality holds by the linearity of $f$ and the final equality holds since $\sigma$ acts as multiplication by $1\in k$ on $\mathbbm{1}$.
For $\sigma\in J_{m}\cap P_{m}$, $\left(\varphi(f)(x)\right)\sigma=0$ since $\sigma$ acts as multiplication by $0\in k$ on $\mathbbm{1}$. On the other hand, $\varphi(f)(x\sigma)=f([x\sigma])=f([0])=0$ since $\sigma\in J_{m}$ and so $[x\sigma]=[0]\in P_{n}/J_{m}$. Therefore the map $\varphi$ is well-defined.
To show that the map $\psi$ is well-defined, we must show that for $x\in P_{n}$ and $\sigma\in P_{m}$, $\psi(g)([x\sigma])=\psi(g)([x])\sigma$. By the above, it suffices to show this for $\sigma\in\Sigma_{m}$ and $\sigma\in J_{m}\cap P_{m}$.
If $\sigma\in\Sigma_{m}$ then $\psi(g)([x\sigma])=g(x\sigma)=g(x)\sigma=\left(\psi(g)(x)\right)\sigma$ by linearity, and since $\sigma$ acts as multiplication by $1\in k$ on $\mathbbm{1}$.
If $\sigma\in J_{m}\cap P_{m}$, then $\psi(g)([x\sigma])=\psi(g)([0])=g(0)=0$ since $\sigma\in J_{m}$ and so $[x\sigma]=[0]\in P_{n}/J_{m}$. On the other hand, $\left(\psi(g)([x])\right)\sigma=0$ since $\sigma$ acts as multiplication by $0\in k$ on $\mathbbm{1}$. Therefore the map $\psi$ is also well-defined and this completes the proof. ∎
Theorem 5.2.4.
Let $n\geqslant m\geqslant 0$. Suppose that $\delta\in k$ is invertible or that $m<n$. The maps
$\pi^{\ast}\colon\operatorname{Ext}_{k[\Sigma_{n}]}^{\star}\left(\mathbbm{1},\operatorname{Hom}_{k[\Sigma_{m}]}\left(k[\Sigma_{n}],\mathbbm{1}\right)\right)\rightarrow\operatorname{Ext}_{P_{n}}^{\star}\left(\mathbbm{1},\operatorname{Hom}_{P_{m}}(P_{n},\mathbbm{1})\right)$
and
$i^{\ast}\colon\operatorname{Ext}_{P_{n}}^{\star}\left(\mathbbm{1},\operatorname{Hom}_{P_{m}}(P_{n},\mathbbm{1})\right)\rightarrow\operatorname{Ext}_{k[\Sigma_{n}]}^{\star}\left(\mathbbm{1},\operatorname{Hom}_{k[\Sigma_{m}]}\left(k[\Sigma_{n}],\mathbbm{1}\right)\right)$
induced by the inclusion $k[\Sigma_{n}]\rightarrow P_{n}$ and projection $\pi\colon P_{n}\rightarrow k[\Sigma_{n}]$ are mutually inverse isomorphisms.
Proof.
The map
$\sigma\colon\operatorname{Ext}_{k[\Sigma_{n}]}^{\star}\left(\mathbbm{1},\operatorname{Hom}_{k[\Sigma_{m}]}(k[\Sigma_{n}],\mathbbm{1})\right)\rightarrow\operatorname{Ext}_{k[\Sigma_{m}]}^{\star}(\mathbbm{1},\mathbbm{1})$
is an isomorphism by the cohomological version of Shapiro’s lemma.
We have the induced map
$i^{\ast}\colon\operatorname{Ext}_{P_{n}}^{\star}\left(\mathbbm{1},\operatorname{Hom}_{P_{m}}(P_{n},\mathbbm{1})\right)\rightarrow\operatorname{Ext}_{k[\Sigma_{n}]}^{\star}\left(\mathbbm{1},\operatorname{Hom}_{k[\Sigma_{m}]}\left(k[\Sigma_{n}],\mathbbm{1}\right)\right).$
We will construct an isomorphism of graded $k$-modules
$\theta\colon\operatorname{Ext}_{k[\Sigma_{m}]}^{\star}(\mathbbm{1},\mathbbm{1})\rightarrow\operatorname{Ext}_{P_{n}}^{\star}\left(\mathbbm{1},\operatorname{Hom}_{P_{m}}(P_{n},\mathbbm{1})\right)$
and show that the composite $\sigma\circ i^{\ast}\circ\theta$ is the identity on $\operatorname{Ext}_{k[\Sigma_{m}]}^{\star}(\mathbbm{1},\mathbbm{1})$.
We begin by constructing $\theta$. Let $R_{\star}$ be a free resolution of $\mathbbm{1}$ by right $P_{n}$-modules.
The graded $k$-module $\operatorname{Ext}_{P_{n}}^{\star}\left(\mathbbm{1},\operatorname{Hom}_{P_{m}}(P_{n},\mathbbm{1})\right)$ is the cohomology of the cochain complex
$\operatorname{Hom}_{P_{n}}\left(R_{\star},\operatorname{Hom}_{P_{m}}(P_{n},\mathbbm{1})\right).$
By Lemma 5.2.3, there is an isomorphism of cochain complexes
$\operatorname{Hom}_{P_{n}}\left(R_{\star},\operatorname{Hom}_{P_{m}}(P_{n},\mathbbm{1})\right)\cong\operatorname{Hom}_{P_{n}}\left(R_{\star},\operatorname{Hom}_{k[\Sigma_{m}]}(P_{n}/J_{m},\mathbbm{1})\right).$
By the tensor-Hom adjunction, we have an isomorphism of chain complexes
$\operatorname{Hom}_{P_{n}}\left(R_{\star},\operatorname{Hom}_{k[\Sigma_{m}]}(P_{n}/J_{m},\mathbbm{1})\right)\cong\operatorname{Hom}_{k[\Sigma_{m}]}\left(R_{\star}\otimes_{P_{n}}(P_{n}/J_{m}),\mathbbm{1}\right).$
As shown in the proof of [BHP23, Theorem 5.1] (which follows the structure of the proof of [BHP21, Theorem 4.1]), the complex $R_{\star}\otimes_{P_{n}}(P_{n}/J_{m})$ is a free resolution of $\mathbbm{1}$ by right $k[\Sigma_{m}]$-modules, under the assumption that either $\delta$ is invertible or $m<n$. Let $\theta$ be the induced map on cohomology under the isomorphism of chain complexes
$\operatorname{Hom}_{k[\Sigma_{m}]}\left(R_{\star}\otimes_{P_{n}}(P_{n}/J_{m}),\mathbbm{1}\right)\rightarrow\operatorname{Hom}_{P_{n}}\left(R_{\star},\operatorname{Hom}_{P_{m}}(P_{n},\mathbbm{1})\right).$
It remains to show that the composite $\sigma\circ i^{\ast}\circ\theta$ is the identity on $\operatorname{Ext}_{k[\Sigma_{m}]}^{\star}(\mathbbm{1},\mathbbm{1})$.
Following the argument in the proof of [BHP21, Theorem 4.1], we take a projective resolution, $Q_{\star}$, of $\mathbbm{1}$ by right $k[\Sigma_{n}]$-modules and choose a chain map $\tilde{i}\colon Q_{\star}\rightarrow R_{\star}$ that lies over the identity map on $\mathbbm{1}$ and respects the inclusion $k[\Sigma_{n}]\rightarrow P_{n}$. The map
$Q_{\star}\rightarrow R_{\star}\otimes_{P_{n}}(P_{n}/J_{m})$
determined by $q\mapsto\tilde{i}(q)\otimes(1+J_{m})$ is a map of projective $k[\Sigma_{m}]$-resolutions of $\mathbbm{1}$, lying above the identity map on $\mathbbm{1}$ and respecting the module structure. It is therefore a chain homotopy equivalence.
The composite $\sigma\circ i^{\ast}\circ\theta$ is obtained by applying the functor $\operatorname{Hom}_{P_{m}}(\mathbbm{1},-)$ to this chain homotopy equivalence and taking cohomology. ∎
Corollary 5.2.5.
Suppose $\delta\in k$ is invertible. There is an isomorphism of graded $k$-modules
$\operatorname{Ext}_{P_{n}(\delta)}^{\star}\left(\mathbbm{1},\mathbbm{1}\right)\cong H^{\star}\left(\Sigma_{n},\mathbbm{1}\right).$
Proof.
Since $\delta$ is invertible, we can apply Theorem 5.2.4 with $m=n$ and use the identifications $\operatorname{Hom}_{P_{n}}\left(P_{n},\mathbbm{1}\right)\cong\mathbbm{1}$ and $\operatorname{Hom}_{k[\Sigma_{n}]}\left(k[\Sigma_{n}],\mathbbm{1}\right)\cong\mathbbm{1}$. ∎
6. (Co)homology of Tanabe algebras, totally propagating partition algebras and the uniform block algebras
In this section, we will show that for any parameter $\delta$ and any $r\geqslant 2$, the (co)homology of the Tanabe algebra $\mathcal{T}_{n}(\delta,r)$ is isomorphic to the (co)homology of the symmetric group $\Sigma_{n}$. We also prove that the (co)homology of the totally propagating partition algebras and the (co)homology of the uniform permutation algebras are isomorphic to the (co)homology of the symmetric groups. The argument is similar in all three cases.
Definition 6.0.1.
Let $\delta\in k$. Let $r\geqslant 2$. For distinct $i$ and $j$ in $\underline{n}$ with $i<j$, we let $L_{i,j}$ denote the left ideal in $\mathcal{T}_{n}(\delta,r)$ (resp. $TPP_{n}$, $U_{n}$) spanned $k$-linearly by the diagrams where $\overline{i}$ and $\overline{j}$ are in the same connected component.
Lemma 6.0.2.
Let $r\geqslant 2$. Recall that $I_{n-1}$ is the two-sided ideal in $\mathcal{T}_{n}(\delta,r)$ (resp. $TPP_{n}$, $U_{n}$) spanned $k$-linearly by $n$-diagrams having fewer than $n$ propagating components. The left ideals $L_{i,j}$ of Definition 6.0.1 form a $k$-free idempotent left cover of $I_{n-1}$ such that the height is equal to the width. In other words, for each algebra
  1. (1)
    the $L_{i,j}$ cover $I_{n-1}$ and
  2. (2)
    any intersection of the $L_{i,j}$ is either zero or a principal left ideal generated by an idempotent.
Proof.
Recall from Subsection 2.2 that $\kappa(C)$ is the absolute value of the difference between the number of vertices in the left-hand column and the number of vertices in the right-hand column of a connected component $C$ in a partition $n$-diagram.
  1. (1)
    Firstly, we must show that the left ideals $L_{i,j}$ cover $I_{n-1}$ in each case.
    If a partition lies in $L_{i,j}$, it contains a connected component with at least two vertices in the right-hand column and so can have at most $n-1$ propagating components. Therefore, the partition lies in $I_{n-1}$. This is true for all three cases.
    A diagram in $I_{n-1}\subset\mathcal{T}_{n}(\delta,r)$ contains at most $n-1$ propagating components and each connected component must satisfy $\kappa(C)\equiv 0\,(\text{mod}\,r)$. This means the diagram cannot be a permutation diagram and, since $r\geqslant 2$, we cannot have isolated vertices. It follows that at least one connected component contains (at least) two vertices in the right-hand column and so the diagram lies in some $L_{i,j}$. Hence the $L_{i,j}$ cover $I_{n-1}$ for the Tanabe algebra $\mathcal{T}_{n}(\delta,r)$.
    A diagram in $I_{n-1}\subset U_{n}$ contains at most $n-1$ propagating components and each connected component must have an equal number of left and right vertices. It follows that at least one connected component contains (at least) two vertices in the right-hand column and so the diagram lies in some $L_{i,j}$. Hence the $L_{i,j}$ cover $I_{n-1}$ in $U_{n}$.
    A diagram in $I_{n-1}\subset TPP_{n}$ contains at most $n-1$ propagating components and each connected component must contain both left and right vertices. It follows that at least one connected component contains two vertices in the right-hand column (if each vertex in the right-hand column was in a different connected component, we would have $n$ connected components, by the totally propagating condition). Therefore, the diagram lies in some $L_{i,j}$ and the $L_{i,j}$ cover $I_{n-1}$ in $TPP_{n}$.
  2. (2)
    We break up the proof of the second part of lemma into three steps.
    1. (a)
      This part of the argument follows [Boy24b, Lemma 3.8]. Let $\underline{n}_{<}^{2}$ be the set of indices $(i,j)$ with $1\leqslant i<j\leqslant n$. Let $T\subset\underline{n}_{<}^{2}$. Let
      $J=\bigcap_{(i,j)\in T}L_{i,j}.$
      Take $(a,b)\in\underline{n}_{<}^{2}\setminus T$. Let $\nu_{a,b}$ be the diagram whose connected components are $\left\{a,b,\overline{a},\overline{b}\right\}$ and $\left\{i,\overline{i}\right\}$ for $i\in\underline{n}\setminus\left\{a,b\right\}$. We note that $\nu_{a,b}$ satisfies $\kappa(C)\equiv 0\,(\text{mod}\,r)$ for each $r\geqslant 2$ and that $\nu_{a,b}$ is a totally propagating partition diagram and a uniform block permutation diagram, so the subsequent applies equally well to all three cases. We give the argument for $\mathcal{T}_{n}(\delta,r)$.
      We claim that $J\cdot\nu_{a,b}\subset L_{a,b}\cap J$.
      Let $\rho$ be a diagram in $J$. We must show that $\rho\nu_{a,b}\in L_{a,b}\cap J$. Since $\rho\in J$, for each $(i,j)\in T$, the vertices $\overline{i}$ and $\overline{j}$ are connected in $\rho$. Since the vertices $i$ and $\overline{i}$ form a connected component of $\nu_{a,b}$, as do $j$ and $\overline{j}$, we see that $\overline{i}$ and $\overline{j}$ are connected in the composite $\rho\nu_{a,b}$. Therefore $\rho\nu_{a,b}\in J$. Since $L_{a,b}$ is a left ideal and $\nu_{a,b}\in L_{a,b}$, the composite $\rho\nu_{a,b}\in L_{a,b}$.
    2. (b)
      This part follows [Boy24b, Lemma 3.9]. We observe that right multiplication by $\nu_{a,b}$ gives a retraction of the inclusion map $L_{a,b}\cap J\rightarrow J$. As noted in [Boy24b, Lemma 3.9], right multiplication by $\nu_{a,b}$ acts on an $n$-diagram, $d$, by merging the component of $d$ containing $\overline{a}$ with the component containing $\overline{b}$, whilst preserving all other connected components of $d$. For any $\rho\in L_{a,b}$, $\overline{a}$ and $\overline{b}$ already lie in the same connected component and so $\rho\nu_{a,b}=\rho$.
    3. (c)
      By repeatedly applying the argument of the previous two points, we see that the composite
      $J\hookrightarrow\mathcal{T}_{n}(\delta,r)\xrightarrow{\Pi}J,$
      where $\Pi$ is right multiplication by the product of all $\nu_{i,j}$ for $(i,j)\in T$, is the identity map. It then follows that $J$ is principal and generated by an idempotent by [Boy24b, Lemma 2.6].
Therefore, the family of left ideals $L_{i,j}$ forms an idempotent left cover of $I_{n-1}$ whose height is equal to its width. ∎
We now prove our main theorem, showing that the (co)homology of the Tanabe algebras, for $r\geqslant 2$, coincides with the (co)homology of the symmetric groups, independently of the parameter $\delta$ and the parity of the index $n$. We also include the proof for the totally propagating partition algebras and the uniform block permutation algebras at the same time.
Theorem 6.0.3.
Let $\delta\in k$. Let $r\geqslant 2$. There exist isomorphisms of graded $k$-modules
  1. (1)
    $\operatorname{Tor}_{\star}^{\mathcal{T}_{n}(\delta,r)}\left(\mathbbm{1},\mathbbm{1}\right)\cong H_{\star}(\Sigma_{n},\mathbbm{1})$ and $\operatorname{Ext}_{\mathcal{T}_{n}(\delta,r)}^{\star}\left(\mathbbm{1},\mathbbm{1}\right)\cong H^{\star}(\Sigma_{n},\mathbbm{1})$;
  2. (2)
    $\operatorname{Tor}_{\star}^{TPP_{n}}\left(\mathbbm{1},\mathbbm{1}\right)\cong H_{\star}(\Sigma_{n},\mathbbm{1})$ and $\operatorname{Ext}_{TPP_{n}}^{\star}\left(\mathbbm{1},\mathbbm{1}\right)\cong H^{\star}(\Sigma_{n},\mathbbm{1})$;
  3. (3)
    $\operatorname{Tor}_{\star}^{U_{n}}\left(\mathbbm{1},\mathbbm{1}\right)\cong H_{\star}(\Sigma_{n},\mathbbm{1})$ and $\operatorname{Ext}_{U_{n}}^{\star}\left(\mathbbm{1},\mathbbm{1}\right)\cong H^{\star}(\Sigma_{n},\mathbbm{1})$.
Proof.
We observe that we have isomorphisms of $k$-algebras $\mathcal{T}_{n}(\delta)/I_{n-1}\cong k[\Sigma_{n}]$, $TPP_{n}/I_{n-1}\cong k[\Sigma_{n}]$ and $U_{n}/I_{n-1}\cong k[\Sigma_{n}]$. By Lemma 6.0.2, the two-sided ideal $I_{n-1}$ has a $k$-free idempotent left cover with height equal to the width in all three cases. The homological statements now follow from [Boy24b, Theorem 1.7] and the cohomological statements follow from Proposition 4.2.2. ∎